Tonya's budget constraint for gasoline (G) and clothes (C) each month can be expressed with the equation G = 100 – 2C. If the price of clothing is $10, which of the following consumption bundles lies INSIDE Tonya's budget line?
0 units of clothing and 100 units of gasoline
50 units of clothing and 100 units of gasoline
10 units of clothing and 5 units of gasoline
50 units of clothing and 0 units of gasoline
Answer #1

Answer
Option 3
10 units of clothing and 5 units of gasoline


the budget equation is
G=100-2C
C=0
G=100-2*0=100


the first option is on the line
G=100-2*50=0

the second option is above budget line


G=100-10*2=80 but required 5 units
so the bundle lies inside of it.

G=100-2*50=0
lies on the line